A flagpole is bent over by the wind so the top is now touching the ground 12 feet from the base. The angle the pole makes with the ground is 85 degrees. How tall was the pole before the windstorm? Can you help me solve this and show me how you did it?
Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

---
cosA = adj/hyp
cos(85) = 12/hyp
hyp = 12/cos(85)
hyp = 137.684558948037960
---
hyp = sqrt( bb + hh )
hyp^2 = bb + hh
hh = hyp^2 - bb
h = sqrt( hyp^2 - bb )
h = sqrt( 137.684558948037960^2 - 12^2 )
h = 138.206504089770448
---
original pole height:
hyp + h = 137.684558948037960 + 138.206504089770448
---
answer:
about 275.9 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by stanbon(75887)

A flagpole is bent over by the wind so the top is now touching the ground 12 feet from the base. The angle the pole makes with the ground is 85 degrees. How tall was the pole before the windstorm?
-----
Draw the picture::
You have a right triangle with base = 12 ft
and base angle = 85 degrees.
---
Find the height and the hypotenuse::
---
height = 12*tan(85) = 137.16 ft
hypotenuse = 12/cos(85) = 137.68 ft
-----
Ans: 136.16 + 137.68 = 274.84 ft
------------------------------------
Cheers,
Stan H.
================